OpenStudy (anonymous):

In the given right triangle, find the missing length to the nearest tenth. A. 20.2 ft B. 7.5 ft C. 11.7 ft D. 17.3 ft

OpenStudy (anonymous):



OpenStudy (anonymous):

17 ft*

OpenStudy (anonymous):

A. 20.2 ft

OpenStudy (anonymous):

a^2 + b^2 = c^2 11^2 + 17^2 = c^2 121 + 289 = c^2 sqrt 410 = 20.2ft

OpenStudy (anonymous):

a.